### v2

# Reconstructive Visual Instruction Tuning

Abstract
This paper introduces reconstructive visual instruction tuning (Ross), a family of Large Multimodal Models (LMMs) that exploit vision-centric supervision signals. In contrast to conventional visual instruction tuning approaches that exclusively supervise text outputs, Ross prompts LMMs to supervise visual outputs via reconstructing input images. By doing so, it capitalizes on the inherent richness and detail present within input images themselves, which are often lost in pure text supervision. However, producing meaningful feedback from natural images is challenging due to the heavy spatial redundancy of visual signals. To address this issue, Ross employs a denoising objective to reconstruct latent representations of input images, avoiding directly regressing exact raw RGB values. This intrinsic activation design inherently encourages LMMs to maintain image detail, thereby enhancing their fine-grained comprehension capabilities and reducing hallucinations. Empirically, Ross consistently brings significant improvements across different visual encoders and language models. In comparison with extrinsic assistance state-of-the-art alternatives that aggregate multiple visual experts, Ross delivers competitive performance with a single SigLIP visual encoder, demonstrating the efficacy of our vision-centric supervision tailored for visual outputs.
1 Introduction
The success of GPT-style Large Language Models (LLMs) (Radford et al., 2018; 2019; Brown et al., 2020; OpenAI, 2023b; Yang et al., 2024a; Touvron et al., 2023; Chiang et al., 2023; Dubey et al., 2024) has motivated researchers to adapt LLMs to understand multimodal inputs (Liu et al., 2023a; 2024a; Dai et al., 2023; Bai et al., 2023). Notably, visual instruction tuning approaches (Liu et al., 2023a) demonstrate superior performance with cost-efficient training recipes. Some approaches (Chen et al., 2024b; Li et al., 2024c) even surpass GPT-4V(ision) (OpenAI, 2023a) on benchmark evaluations.
Typically, these Large Multimodal Models (LMMs) based on visual instruction tuning adopt a plug-in architecture, as depicted in Figure 1a, where pre-trained vision-language foundation models such as CLIP (Radford et al., 2021) are responsible for projecting images into visual tokens. They serve as prefix tokens for multimodal comprehension. However, this type of design, i.e., visual encoder connector LLM language instructions, where “” indicates supervision, is primarily LLM-centric: (i) visual comprehension largely depends on vision-to-text alignment and the selected vision models, and (ii) supervision derives exclusively from text data. As a result, they exhibit systematic visual shortcomings such as recognizing specific visual patterns (Tong et al., 2024b).
Until very recently, some concurrent works proposed vision-centric solutions (Tong et al., 2024a; b). Illustrated in Figure 1b, their solutions leverage extrinsic assistance via aggregating several different visual experts. Inspired by the evolution in image recognition, from manually designed visual features (Sánchez & Perronnin, 2011) to learnable deep convolutional models (Krizhevsky et al., 2012), we suggest that intrinsic activation offers a more viable path forward. Just as deep models automatically learn hierarchical and abstract features from raw data, we believe intrinsic activation methods are similarly more adaptable for multimodal comprehension, reducing reliance on hand-crafted engineering, thereby enhancing both generalization and performance. Therefore, we aim to explore intrinsic activation solutions based on the following principles:
1.
Supervise Visual Outputs. Current LMMs solely supervise text outputs, neglecting a significant amount of visual outputs unused. For instance, LLaVA-v1.5 (Liu et al., 2024a) utilizes visual tokens to represent a single image, yet their corresponding outputs remain unsupervised. Intuitively, since input images themselves inherently provide rich and detailed information, we regard LMMs reconstructing input images as the supervision of those visual outputs. This approach encourages LMMs to maintain low-level details, thereby enhancing their fine-grained comprehension abilities and reducing hallucinations.
2.
Explore the Optimal Formulation. Designing this self-supervised task effectively is not straightforward. Motivated by the success of masked autoencoder (He et al., 2022) compared to its basic version denoising autoencoder (Vincent et al., 2008), we identify handling heavy spatial redundancy of visual signals as the underlying key factor. To this end, we formulate our approach as follows: (i) for reconstruction targets, instead of raw RGB pixels, we make LMMs reconstruct latent visual tokens, and (ii) for reconstruction objectives, to avoid directly regressing exact token values, we adopt per-token denoising.
To this end, we propose Ross, termed of reconstructive visual instruction tuning, which utilizes input images as direct supervision signals illustrated in Figure 1c. Technically, to address the spatial redundancy inherent in natural visual signals (He et al., 2022), we train a small denoising network, which takes high-level visual outputs as conditions to recover low-level fine-grained visual tokens , representing an underlying distribution . These latent tokens are derived from a frozen teacher tokenizer such as continuous VAE (Kingma, 2013) and discrete VQGAN (Esser et al., 2021). Unlike extrinsic assistance solutions (Tong et al., 2024a; b), our intrinsic activation solution naturally maintains a lightweight inference procedure. More importantly, when adapting to new visual domains, our solution avoids a careful choice of new domain-specific experts, e.g., MiDaS-3.0 (Birkl et al., 2023) for understanding depth maps, which is more efficient and easier to implement.
Empirically, Ross achieves top performance across a wide range of multimodal comprehension benchmarks. Notably, our Ross excels in fine-grained vision-centric benchmarks (Tong et al., 2024b; Masry et al., 2022) and hallucination benchmarks (Guan et al., 2024; Li et al., 2023c). To be specific, with a single SigLIP (Zhai et al., 2023) as the visual encoder, Ross-7B achieves 57.3 on HallusionBench (Guan et al., 2024) and 54.7 on MMVP (Tong et al., 2024b), significantly outperforms state-of-the-art alternatives with similar model sizes which aggregate several visual experts as extrinsic assistance, e.g., Cambrian-1-8B (Tong et al., 2024a). In-depth analysis demonstrates the effectiveness of Ross for directing focus towards visual elements and understanding depth maps. We hope our research will inspire future work in designing supervision signals for large multimodal models.
2 Related Work
Visual Instruction Tuning. Most visual instruction tuning-based LMMs adopt a plug-in architecture (Liu et al., 2023a; 2024a; Bai et al., 2023; Wang et al., 2024a), where a language-supervised visual encoder (Radford et al., 2021; Zhai et al., 2023) is responsible for extracting visual tokens. A connector is used to map those visual representations into the LLM space, e.g., Resamplers (Alayrac et al., 2022), Q-Formers (Li et al., 2023b; Dai et al., 2023; Bai et al., 2023; Ge et al., 2024a), and MLPs (Liu et al., 2023a; 2024a; Li et al., 2024c; Liu et al., 2024b; Li et al., 2024a). These LMMs usually follow a two-stage training recipe. During the alignment stage, the connector is trained on high-quality caption data. Next, the full model is trained on single-image visual instruction tuning data. However, only text outputs are supervised. Ross, on the other hand, introduces novel vision-centric supervision via reconstructing fine-grained visual tokens conditioned on visual outputs.
Visual Encoders for LMMs. As the original CLIP (Radford et al., 2021) adopted by conventional visual instruction tuning approaches is trained on noisy image-text pairs, it exhibits specific visual shortcomings, and thus stronger backbones (Fang et al., 2024; Zhai et al., 2023; Chen et al., 2024c) have been introduced to LMMs. Some concurrent works (Tong et al., 2024b; a) leverage extrinsic assistance, which further utilizes vision-only self-supervised models (Oquab et al., 2023; Wang et al., 2023a; b; c; He et al., 2022; Caron et al., 2021) and domain experts (Kirillov et al., 2023; Birkl et al., 2023; Rombach et al., 2022). Ross, from a new intrinsic activation perspective, aims to catalyze enhanced comprehension through reconstructing input images with no extra visual experts.
Generative Objectives for LMMs. Another line of work introduces pre-trained text-to-image diffusion models (Rombach et al., 2022) to make LMMs capable of both comprehension and generation (Dong et al., 2024; Ge et al., 2024a; Sun et al., 2024b; Ge et al., 2024b; Sun et al., 2023). Our Ross, with a totally different motivation, targets to catalyze multimodal comprehension via reconstruction. Specifically, conditions are different, where Dong et al. (2024) and Sun et al. (2024b) take outputs corresponding to learnable queries as conditions, while our Ross takes outputs corresponding to visual inputs. Those methods are generative while Ross is reconstructive. The detailed pipeline comparison can be found in Appendix C.
3 Preliminaries
Large Multimodal Models. In the literature (Radford et al., 2018; 2019), a -parameterized LLM models the canonical causal distribution of each text token as , where represents a sequence of text tokens. To make LLMs understand visual contents, typical plug-in style LMMs (Liu et al., 2023a; 2024a) regard a sequence of visual tokens as prefix tokens. Specifically, an input image is first projected into a sequence of visual tokens by a -parameterized visual encoder such as CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), where indicates the spatial resolution. Then, a -parameterized multimodal projector is utilized to project these visual tokens into the feature space of LLMs. As a result, the canonical causal distribution in a multimodal sentence containing an image becomes
where is the parameters and indicates the projected visual tokens. is the number of visual tokens and indicates the feature channel. The visual encoder could be either frozen (Liu et al., 2023a; 2024a; Tong et al., 2024a) or fine-tuned (Liu et al., 2024b; Bai et al., 2023; Li et al., 2024c; Wang et al., 2024c).
Training Recipes for LMMs. LMMs almost follow a two-stage training recipe (Liu et al., 2023a), i.e., the pre-training stage (or the alignment stage) and the supervised fine-tuning stage (or the instruction tuning stage). The instruction (supervision) comes from languages such as the answers to VQA tasks, maximizing the log-likelihood of text outputs:
where represents the number of visual tokens and visual outputs (one input token corresponds to one visual output). From Equation 2, we can tell that only text outputs are supervised.
4 Ross: Reconstructive Visual Instruction Tuning
In this section, we first provide an overview of our reconstructive visual instruction tuning (Ross). Then, we discuss our explorations towards the optimal formulation in the following subsections, with the ultimate goal of handling spatial redundancy of visual signals to provide meaningful visual supervision. Our explorations mainly include reconstruction targets and the training objective.
Overview. Illustrated in Figure 2, the overall philosophy of our Ross is to construct reconstructive visual supervision signals on visual outputs . The training objective includes (i) the original next-token prediction on shown in the right part of Figure 2, and (ii) another reconstructive term in the left part of Figure 2, i.e., . Specifically, this visual term could be any custom measurements between and specific reconstruction targets of image :
where indicates the -parameterized post projection that maps the dimensions of visual tokens to be consistent with the teacher tokenizer .
Variants of Ross. Evidently, different choices of and contribute to different variants. controls the reconstruction target while defines the objective:
1.
Towards the target, can be the pachify operation (Dosovitskiy et al., 2021), resulting in pixel-level reconstruction, or pre-trained fine-grained visual tokenizers such as VAE (Kingma, 2013) and VQGAN (Esser et al., 2021), leading to latent-level reconstruction. could even be vision-only models such as DINOv2 (Oquab et al., 2023), making LMMs learn specific visual patterns from , which is also a type of latent-level reconstruction.
2.
Towards the objective, the most straightforward choice of is MSE or cosine similarity for regressing raw pixel values or latent features, respectively. We also explore the denoising objective (Ho et al., 2020) to avoid being overwhelmed by fitting exact values.
We introduce our explorations step by step in the following sections. The ultimate goal of our exploration is to design an appropriate self-supervised reconstructive pre-text task that provides meaningful vision-centric supervision signals to LMMs, where handling the spatial redundancy of visual signals (He et al., 2022) becomes the crux.
4.1 Ross: Regressing as Reconstructive Visual Instruction
In this section, we introduce straightforward variants, i.e., regressing as reconstructive visual instruction. As shown in Figure 3, depending on the choice of , it mainly has three variants: (a) Ross-Pixel, (b) Ross-Latent, and (c) Ross-Latent2Pixel.
Directly Regressing Raw RGB Values. The most straightforward variant is to directly regress raw RGB values illustrated in Figure 3a, called “Ross-Pixel”. Under such a setting, is the patchify operation (Dosovitskiy et al., 2021), reshaping the image into a sequence of flattened 2D patches , where is the resolution of each image patch and indicates the resulting number of patches. can be a simple MLP, mapping the dimension of visual outputs from to . The measurement is MSE. However, as visual signals suffer from heavy spatial redundancy (He et al., 2022), such a design may not provide meaningful supervision to LMMs. An intuitive alternative to avoid directly regressing raw RGB values while still reconstructing the image is to urge LMMs to reconstruct latent tokens, introduced as follows.
Regressing Latent Tokens. Illustrated in Figure 3b, Ross-Latent aims to regress fine-grained latent tokens extracted by the teacher tokenizer. can be models trained with discriminative tasks such as DINOv2 (Oquab et al., 2023) and DEIT-III (Touvron et al., 2022). The encoder part of models trained with reconstruction tasks such as VQGAN (Esser et al., 2021) and VAE (Kingma, 2013) are also capable. here is the consine-similarity. Intuitively, the decoder part of the latter is able to remap latent tokens into the pixel space. Therefore, supervising in the pixel space via decoding becomes another valid variant introduced as follows.
Regressing RGB Values via Decoding. Shown in Figure 3c, Ross-Latent2Pixel requires a decoder to project predicted latent tokens into the RGB pixel space, resulting in predicted image . Let be the decoder part of VQGAN (Esser et al., 2021) or VAE (Kingma, 2013), and the regressive MSE objective is performed on pixel-space. Note that we simply use to represent the decoding process, which is actually not the inverse function of mathematically.
Discussion. Recall that we need to find the optimal solution to address the spatial redundancy of natural visual signals, the target-level exploration above achieves this goal partially, as the objective is limited to vanilla regression. To this end, inspired by Ho et al. (2020) and Li et al. (2024e), we further incorporate a novel denoising objective in the following section.
4.2 Ross: Denoising as Reconstructive Visual Instruction
As an objective for handling heavy spatial redundancy to provide meaningful vision-centric supervision signals, denoising is better than vanilla regressing, since the introduction of noise into the training data acts as an implicit form of data augmentation and regularization. The denoising process encourages the model to focus on the underlying data manifold rather than memorizing specific instance values (Chen et al., 2023c; Song & Ermon, 2019; Karras et al., 2022; Yang et al., 2024b).
Techinically, as illustrated in Figure 4a, our final Ross takes high-level visual outputs as conditions to recover clean fine-grained tokens from noisy tokens . Specifically, clean tokens are obtained from the teacher tokenizer . By default, we utilize a continuous VAE (Kingma, 2013) regularized by Kullback–Leibler (KL) divergence provided by Rombach et al. (2022), since it is believed to capture sufficient image details. The training procedure of the denoiser follows a diffusion process (Ho et al., 2020):
The denoiser actually estimates the conditional expectation . More details about the background knowledge of diffusion models can be found in Appendix A.
Architecture of the Denoiser. As conditions are causal, we introduce a self-attention module to model the inter-token dependencies illustrated in Figure 4b. Specifically, the architecture of the denoiser is a stack of Transformer Encoder blocks (Vaswani et al., 2017) and each block contains three extra projections for conditions , inputs , and timesteps , respectively.
Choices of the Teacher Tokenizer. By default, we adopt latent denoising and we take a continuous tokenizer provided by Rombach et al. (2022) as , since it manages to reconstruct input images with a low rFID (Heusel et al., 2017) and thus it is expected to preserve many low-level details of input images. This extra reconstructive objective, however, is not limited to any certain tokenizer . Discrete tokenizers such as VQGAN (Esser et al., 2021), and vision self-supervised models such as DINOv2 (Oquab et al., 2023), are also qualified to be the tokenizer. Even the patchify operation (Dosovitskiy et al., 2021) is capable, resulting in pixel denoising.
5 Experiments
5.1 Ablation Study
Implementation Details. All ablation studies are implemented based on LLaVA-v1.5 (Liu et al., 2024a). The visual encoder is CLIP-ViT-L/14@336 (Radford et al., 2021) and the base LLM is Qwen2-7B-Instruct (Yang et al., 2024a). The training data is LLaVA-558K (Liu et al., 2023a) and Cambrian-737K (Tong et al., 2024a) for the pre-training stage and the instruction tuning stage, respectively. We evaluate our each variant of Ross mainly on (i) hallucination: POPE (Li et al., 2023c) and HallusionBench (Guan et al., 2024), (ii) fine-grained comprehension: MMVP (Tong et al., 2024b) and ChartQA (Masry et al., 2022), and (iii) general comprehension: MMBench (Liu et al., 2023b) English dev split. All evaluations are conducted with VLMEvalKit (Duan et al., 2024). Evaluation prompts can be found in Appendix B.
Pixel Regression v.s. Latent Regression. Starting from the visual instruction tuning baseline (Liu et al., 2023a; 2024a), we first explore the effectiveness of using regression as the objective for our reconstructive visual instruction tuning. We utilize a continuous VAE (Kingma, 2013) with an encoder-decoder architecture provided by Rombach et al. (2022), where the encoder part serves as for Ross-Latent while the decoder part is for Ross-Latent2Pixel. As illustrated in Figure 8, our vision-centric regression supervision outperforms the visual instruction tuning baseline in most cases. Moreover, latent regression performs the best since regressing raw RGB pixels fails to provide meaningful supervision signals, regardless of whether utilizing a decoder or not.
Choices of . We study the effectiveness across different latent teacher tokenizers in Figure 8, including KL-16 provided by Rombach et al. (2022), which is a continuous VAE (Kingma, 2013) with Kullback–Leibler (KL) divergence, self-supervised DINOv2 (Oquab et al., 2023), fully-supervised DEiT-III (Touvron et al., 2022), and language-supervised EVA02CLIP (Fang et al., 2024). Among them, KL-16 is the best choice. One intuitive explanation is that it is expected to preserve the most image details, since it was originally designed to accurately reconstruct input images.
Regression v.s. Denoising. In Figure 8, we study the effectiveness of the denoising objective over vanilla regression across different tokenizers, i.e., KL-16 (Rombach et al., 2022) and DINOv2 (Oquab et al., 2023). Notably, even if Ross-Latent (KL-16) has already outperformed the visual instruction tuning baseline by a large margin, Ross manages to bring significant improvements by replacing regression with denoising. Therefore, denoising is better at handling visual spatial redundancy.
Finally, we leverage the insights and conclusions from all our previous studies to train our Ross. Specifically, we regard the optimal formulation Ross (KL-16), i.e., denoising with the KL-16 tokenizer, as our final Ross. Please refer to Section C.1 for ablations on the architecture of the denoiser, continuous tokenizer v.s. discrete tokenizer, and the denoising schedule.
5.2 In-Depth Analysis
Attention Analysis. We compute the attention scores of the last token with respect to all visual tokens on MMVP (Tong et al., 2024b). Quantitative and qualitative comparisons between the visual instruction tuning baseline (LLaVA) (Liu et al., 2024a) and our Ross are provided in Table 1 and Figure 8, respectively. Table 1 reveals that the attention scores achieved by Ross are significantly higher than those of LLaVA, indicating that the inclusion of vision-centric reconstructive objective effectively directs focus towards input images, thereby enhancing the comprehending visual signals. Similarly, Figure 8 demonstrate that the implementation of enables the alignment of attention closely with the relevant visual elements corresponding to the text query.
| Statistic (10) | LLaVA | Ross | P-value |
| Mean | 2.03 | 2.36 | 1.2710 |
| 25th Percentile | 1.50 | 1.81 | – |
| Median | 1.90 | 2.26 | 4.3910 |
| 75th Percentile | 2.42 | 2.76 | – |
| 95th Percentile | 3.51 | 3.69 | – |
| Method | SFT Data | w/ | Hallucination | Fine-Grained | General | |||
| 737K | 102K | POPE | Hallu. | MMVP | ChartQA | MMB | ||
| Baseline | 737K + Caption 102K | – | – | 86.2 | 55.1 | 32.0 | 30.9 | 74.4 |
| Reconstructive | 737K + Caption 102K | ✓ | ✓ | 87.6 | 58.0 | 38.7 | 40.4 | 75.2 |
| Reconstructive | 737K + Caption 102K | – | ✓ | 87.6 | 56.3 | 37.3 | 39.7 | 74.9 |
| Generative | 737K + Creation 102K | – | ✓ | 85.4 | 52.0 | 30.0 | 31.2 | 73.9 |
| Language Model | POPE | Hallu. | MMVP | ChartQA | OCRBench | MMB | |
| Visual Encoder: CLIP-ViT-L/14@336 | |||||||
| Vicuna-7B-v1.5 | – | 86.3 | 52.5 | 28.0 | 32.9 | 339 | 67.0 |
| ✓ | 87.2 0.9 | 55.8 3.3 | 36.0 8.0 | 39.8 6.9 | 350 11 | 67.6 0.6 | |
| Qwen2-7B-Instruct | – | 87.9 | 55.0 | 29.3 | 34.0 | 363 | 73.8 |
| ✓ | 88.4 0.5 | 56.7 1.7 | 42.0 12.7 | 37.1 3.1 | 381 18 | 75.2 1.4 | |
| Visual Encoder: SigLIP-ViT-SO400M/14@384 | |||||||
| Vicuna-7B-v1.5 | – | 86.0 | 50.4 | 27.3 | 36.2 | 354 | 64.5 |
| ✓ | 86.8 0.8 | 53.2 2.8 | 38.0 10.7 | 41.6 5.4 | 365 11 | 65.7 1.2 | |
| Qwen2-7B-Instruct | – | 88.5 | 57.3 | 40.7 | 44.4 | 432 | 76.3 |
| ✓ | 88.7 0.2 | 58.2 0.9 | 49.3 8.6 | 46.3 1.9 | 448 16 | 76.9 0.6 | |
Generative v.s. Reconstructive. We ablate the effectiveness of reconstruction over generation in Table 3. Similar to Sun et al. (2024b) and Dong et al. (2024), for generative cases, we adopt 576 learnable latent tokens to query the LMM and utilize the corresponding outputs as conditions to the denoiser. The detailed pipeline of these two methods can be found at Figure 11 in Section C.1. However, generative methods require specific creation data and can not be naively implemented on the original SFT data. To build creation data, we utilize GPT-4o to transfer 102K caption into text-to-image creation data from ShareGPT4V (Chen et al., 2023b) and combine them with the original SFT data. From Table 3, we can tell that reconstructive objectives boost comprehension while generative alternatives cannot. An intuitive explanation of this evidence can be found in Section C.1.
Ross with Different LLMs and Visual Encoders. To demonstrate the effectiveness of our vision-centric supervision adopted by our Ross, we conduct systematic experiments across different base LLMs and visual encoders. From Table 3, Ross contributes to significant improvements consistently, especially on fine-grained comprehension benchmarks, i.e., MMVP (Tong et al., 2024b) and ChartQA (Masry et al., 2022). Extended experiments on more representative benchmarks can be found at Table 12 in Section C.1.
| Model | POPE | Hallu. | MMB | MMB | SEED | MMMU | MMVP | GQA | AI2D |
| GPT-4V-1106 (OpenAI, 2023a) | 75.4 | 65.8‡ | 75.8 | 75.1‡ | 71.6 | 53.8 | 50.0 | 36.8 | 78.2 |
| Gemini-1.5 Pro (Team et al., 2023) | – | – | 73.6 | – | 70.7 | 47.9 | – | – | – |
| MM-1-8B (McKinzie et al., 2024) | 86.6 | – | 72.3 | – | 69.9 | 37.0 | – | 72.6 | – |
| Mini-Gemini-8B (Li et al., 2024f) | – | – | 72.7 | – | 73.2 | 37.3 | 18.7 | 64.5 | 73.5 |
| DeepSeek-VL-7B (Lu et al., 2024) | 85.8‡ | 44.1‡ | 73.2 | 72.8 | 70.4 | 36.6 | – | – | 64.9‡ |
| Cambrian-1-8B (Tong et al., 2024a) | 87.4‡ | 48.7‡ | 75.9 | 68.9‡ | 74.7 | 42.7 | 51.3 | 64.6 | 73.0 |
| Ross-7B | 88.3 | 57.1 | 79.1 | 77.1 | 73.6 | 46.6 | 56.7 | 65.5 | 79.3 |
| Base LLM: Vicuna-7B-v1.5 | |||||||||
| LLaVA-v1.5-7B‡ (Liu et al., 2024a) | 86.2 | 47.5 | 65.5 | 58.5 | 66.0 | 34.4 | 20.0 | 62.0 | 55.4 |
| LLaVA-v1.6-7B‡ (Liu et al., 2024b) | 86.5 | 35.8 | 67.4 | 60.1 | 70.2 | 35.8 | 37.3 | 64.2 | 67.1 |
| Ross-7B | 88.2 | 55.2 | 67.7 | 61.3 | 67.6 | 36.9 | 39.3 | 63.7 | 69.3 |
| Base LLM: Vicuna-13B-v1.5 | |||||||||
| LLaVA-v1.5-13B‡ (Liu et al., 2024a) | 82.5 | 44.9 | 68.8 | 63.6 | 68.2 | 36.6 | 32.0 | 63.3 | 60.8 |
| LLaVA-v1.6-13B‡ (Liu et al., 2024b) | 86.2 | 36.7 | 70.0 | 64.1 | 71.9 | 36.2 | 35.3 | 65.4 | 72.4 |
| Mini-Gemini-13B (Li et al., 2024f) | – | – | 68.6 | – | 73.2 | 37.3 | 19.3 | 63.7 | 70.1 |
| Cambrian-1-13B (Tong et al., 2024a) | 85.7‡ | 54.0‡ | 75.7 | 65.9‡ | 74.4 | 40.0 | 41.3 | 64.3 | 73.6 |
| Ross-13B | 88.7 | 56.4 | 73.6 | 67.4 | 71.1 | 41.3 | 44.7 | 65.2 | 73.8 |
Reconstruction Results. We fine-tune the denoiser to recover latent tokens from a frozen KL-16 provided by Rombach et al. (2022) conditioned on Ross-7B features on ImageNet-1K (Deng et al., 2009) for only five epochs, where the denoiser manages to produce reasonable reconstruction results as illustrated in Figure 9. This interesting finding demonstrates that high-level Ross-7B features actually contain image details. We hope this finding will inspire future work.
Computational Overhead. The denoising process introduces a negligible increase in training time (10% compared to the baseline), while the benefits outweigh the minor additional costs. Please refer to Table 10 in Appendix B for details.
5.3 Comparison with State-of-the-Arts
Ross utlizes a single SigLIP-ViT-SO400M/14@384 (Zhai et al., 2023) as the visual encoder. Ross-7B utilizes Qwen2-7B-Instruct (Yang et al., 2024a) and Ross-13B adopts Vicuna-13B-v1.5 (Chiang et al., 2023) as the base LLM. The implementation almost follows LLaVA-v1.5 (Liu et al., 2024a) without the high-resolution image-slicing technique (Liu et al., 2024b). Thus, our primary comparison of Ross with alternative methods focuses on benchmarks that do not require exceptionally high-resolution inputs. We use a mixture of 2M caption data for the pre-training stage, which consists of 1246K from ShareGPT4V (Chen et al., 2023b) and 707K from ALLaVA (Chen et al., 2024a). The instruction tuning data is a mixture of Cambrian-737K (Tong et al., 2024a) and SMR-473K (Zhang et al., 2024). We further incorporate our Ross with the “anyres” technique (Liu et al., 2024b) and compare with others on high-resolution benchmarks at Table 13 in Section C.2.
Illustrated in Table 4, we compare our Ross with both private models (OpenAI, 2023a; Team et al., 2023; McKinzie et al., 2024) and open-sourced alternatives (Liu et al., 2024a; b; Tong et al., 2024a; Li et al., 2024f; Lu et al., 2024). The previous open-source state-of-the-art Cambrian-1 (Tong et al., 2024a) leverages extrinsic assistance that aggregates CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), DINOv2 (Oquab et al., 2023), and ConvNext (Liu et al., 2022). On the other hand, our Ross stands for intrinsic activation. With only a single SigLIP (Zhai et al., 2023) model as the visual encoder, our Ross surpasses Cambrian-1 (Tong et al., 2024a), under most cases, without a careful choice of the visual experts and naturally maintains a lightweight inference procedure. Ross is also data-efficient compared with Cambrian-1 (Tong et al., 2024a), since it requires 7M instruction tuning data. Notably, Ross-7B even surpasses GPT-4V-1106 and Gemini-1.5 Pro on several benchmarks such as POPE (Li et al., 2023c), MMBench (Liu et al., 2023b), and MMVP (Tong et al., 2024b).
| Method | Test Inputs | MiDaS | Size | Reaching | Position | Existence | Counting | Average | |
| LLaVA | RGB | – | – | 20.0 | 51.7 | 58.8 | 70.0 | 74.6 | 55.0 |
| RGB + D | – | – | 21.7 1.7 | 45.0 6.7 | 58.8 – 0.0 | 65.0 5.0 | 77.7 3.1 | 53.6 1.4 | |
| RGB | – | ✓ | 21.7 | 60.0 | 64.7 | 80.0 | 84.1 | 62.1 | |
| RGB + D | – | ✓ | 21.7 – 0.0 | 51.7 8.3 | 70.6 5.9 | 65.0 15.0 | 91.1 7.0 | 60.0 2.1 | |
| Ross | RGB | ✓ | – | 25.0 | 53.3 | 64.7 | 70.0 | 75.3 | 57.7 |
| RGB + D | ✓ | – | 28.3 3.3 | 65.0 11.7 | 67.6 2.9 | 85.0 15.0 | 84.6 8.7 | 66.1 8.4 | |
| GPT-4o | RGB | – | – | 43.3 | 51.7 | 70.6 | 85.0 | 84.5 | 67.0 |
| RGB + D | – | – | 40.0 3.3 | 51.7 – 0.0 | 61.8 8.8 | 90.0 5.0 | 85.2 0.7 | 65.7 1.3 |
5.4 Applications
Transfer Learning on Understanding Depth Maps. We further evaluate the transfer learning capability of our Ross on SpatialBench (Cai et al., 2024), which requires the model to understand depth maps. We compare our Ross with the visual instruction tuning baseline, with the same training data and model architecture. Also, we compare the effectiveness of the extrinsic assistance solution, i.e., combining a depth expert MiDaS-3.0 (Birkl et al., 2023) to visual instruction tuning, with our intrinsic activation solution. Specifically, the pre-training data is LLaVA-558K (Liu et al., 2023a) and the fine-tuning data is SpatialQA-853K (Cai et al., 2024), where each conversation contains the RGB image and the depth maps extracted by ZoeDepth (Bhat et al., 2023). The visual encoder is CLIP-ViT-L/14@336 (Radford et al., 2021) and the base LLM is Qwen2-7B-Instruct (Yang et al., 2024a). As demonstrated in Table 5, our Ross manages to make use of the extra depth map, as consistent and significant improvements are observed when taking “RGB + D” inputs for testing. Extrinsic assistance approaches cannot take advantage of extra depth maps when testing. Even GPT-4o cannot fully understand depth maps. Qualitative results can be found at Figure 16 in Section C.4.
6 Conclusion
This paper introduces reconstructive visual instruction tuning (Ross), leveraging a vision-centric reconstructive objective to supervise visual outputs. To avoid being overwhelmed by heavily redundant raw RGB values, we train a denoiser to recover clean latent visual representations conditioning on visual outputs. Experimentally, the proposed objective indeed brings enhanced comprehension capabilities and reduced hallucinations. Ross outperforms the state-of-the-art under most cases with only a single SigLIP (Zhai et al., 2023) as the visual encoder. The in-depth analysis demonstrates that high-level features from Ross-7B actually contain sufficient details for low-level image reconstruction. This finding reveals the possibility of equipping comprehension LMMs with the ability of naive generation without the help of generation experts such as Stable Diffusion (Rombach et al., 2022).
Acknowledgements
The work was supported by the National Science and Technology Major Project of China (No. 2023ZD0121300), the National Natural Science Foundation of China (No. U21B2042, No. 62320106010), the 2035 Innovation Program of CAS, and the InnoHK program.
Appendix
Appendix A Latent Diffusion Models
Given a set of clean latent tokens drawn from , the forward diffusion process is a Markov chain that gradually adds random Gaussian noise to the original sample:
where denotes the Gaussian distribution, and indicates discrete timesteps. indicates a pre-defined time-dependent variance schedule. According to Ho et al. (2020), to admit sampling at an arbitrary timestep directly from , this transition can be reformulated as
where and . A latent diffusion model learns to reverse this progressive noise addition process for latent tokens. Specifically, to iteratively generate clean tokens from pure noise conditioned on , we need to reverse the forward process by
where a -parameterized neural network is trained to predict the added noise during the forward process. indicates the posterior noise variance. The training objective of is
Appendix B Implementation Details
| Config | Stage I | Stage II |
| Trainable parts | projector + denoiser | projector + LLM + denoiser |
| Frozen parts | visual encoder + LLM + teacher tokenizer | visual encoder + teacher tokenizer |
| Global batch size | 256 | 128 |
| Batch size per GPU | 16 | 4 |
| Accumulated steps | 2 | 4 |
| DeepSpeed zero stage | 2 | 3 |
| Learning rate | 110 | 210 |
| Learning rate schedule | warmup + cosine decay | |
| Warmup ratio | 0.03 | |
| Weight decay | 0 | |
| Epoch | 1 | |
| Optimizer | AdamW | |
| Precision | bf16 | |
| Dataset | # Samples |
| LLaVA (Liu et al., 2023a) | 158K |
| ShareGPT (Team, 2023) | 40K |
| VQAv2 (Goyal et al., 2017) | 83K |
| GQA (Hudson & Manning, 2019) | 72.1K |
| OKVQA (Marino et al., 2019) | 9K |
| OCRVQA (Mishra et al., 2019) | 80K |
| A-OKVQA (Schwenk et al., 2022) | 50K |
| TextVQA (Singh et al., 2019) | 21.9K |
| RefCOCO (Kazemzadeh et al., 2014) | 30K |
| VG (Krishna et al., 2017) | 86.4K |
| DVQA (Kafle et al., 2018) | 13K |
| DocVQA (Mathew et al., 2021) | 15K |
| ChartQA (Masry et al., 2022) | 28.1K |
| AI2 Diagrams (Kembhavi et al., 2016) | 15.5K |
| Dataset | # Samples |
| ScienccQA (Saikh et al., 2022) | 9K |
| TextbookQA (Kembhavi et al., 2017) | 9.5K |
| AI2 Diagrams (Kembhavi et al., 2016) | 12.4K |
| ChartQA (Masry et al., 2022) | 28.3K |
| DVQA (Kafle et al., 2018) | 200K |
| ArxivQA (Li et al., 2024d) | 100K |
| GeoQA3 (Chen et al., 2021) | 5K |
| Geometry3K (Lu et al., 2021) | 2.1K |
| GeoQA+ (Cao & Xiao, 2022) | 72.3K |
| MathVision (Wang et al., 2024b) | 2.7K |
| TabMWP (Lu et al., 2022) | 30.7K |
| Benchmark | Response formatting prompts |
| POPE (Li et al., 2023c) | – |
| HallusionBench (Guan et al., 2024) | Answer the question using a single word or phrase. |
| MMBench (Liu et al., 2023b) | Answer with the option’s letter from the given choices directly. |
| SEED-Bench (Li et al., 2023a) | Answer with the option’s letter from the given choices directly. |
| MMMU (Yue et al., 2024) | Answer with the option’s letter from the given choices directly. |
| MMVP (Tong et al., 2024b) | Answer with the option’s letter from the given choices directly. |
| AI2D (Hiippala et al., 2021) | Answer with the option’s letter from the given choices directly. |
| RealWorldQA (xAI, 2024) | Answer with the option’s letter from the given choices directly. |
| GQA (Hudson & Manning, 2019) | Answer the question using a single word or phrase. |
| ChartQA (Masry et al., 2022) | Answer the question using a single word or phrase. |
| OCRBench (Liu et al., 2023c) | Answer the question using a single word or phrase. |
| DocVQA (Mathew et al., 2021) | Answer the question using a single word or phrase. |
| InfoVQA (Biten et al., 2022) | Answer the question using a single word or phrase. |
| TextVQA (Singh et al., 2019) | Answer the question using a single word or phrase. |
| Vision | Base LLM |
Speed (s/iter) | Time |
| CLIP-L/336 | Qwen2-7B-Instruct | – | 7.63 B | 8.31 | 13h 17min | 45.34 G | ||||
| CLIP-L/336 | Qwen2-7B-Instruct | ✓ | 7.68 B | 9.02 (1.09) | 14h 25min | 46.62 G (1.03) | ||||
| CLIP-L/336 | Vicuna-13B-v1.5 | – | 13.05 B | 13.33 | 21h 19min | 48.62 G | ||||
| CLIP-L/336 | Vicuna-13B-v1.5 | ✓ | 13.11 B | 14.69 (1.10) | 23h 30min | 49.07 G (1.01) | ||||
| SigLIP-L/384 | Qwen2-7B-Instruct | – | 7.63 B | 8.77 | 14h 1min | 47.08 G | ||||
| SigLIP-L/384 | Qwen2-7B-Instruct | ✓ | 7.68 B | 9.48 (1.08) | 15h 9min | 52.07 G (1.11) | ||||
| SigLIP-L/384 | Vicuna-13B-v1.5 | – | 13.05 B | 14.22 | 22h 44min | 48.80 G | ||||
| SigLIP-L/384 | Vicuna-13B-v1.5 | ✓ | 13.11 B | 15.32 (1.08) | 24h 30min | 52.68 G (1.08) |
Hyperparameters. The hyperparameters of Ross are provided in Table 6. We simply borrow most configurations from LLaVA-v1.5 (Liu et al., 2024a) without further tuning, as we find it works well with our Ross, even if we adopt SigLIP (Zhai et al., 2023) and Qwen2 (Yang et al., 2024a) while the original LLaVA-v1.5 (Liu et al., 2024a) utilized CLIP (Radford et al., 2021) and Vicuna-v1.5 (Chiang et al., 2023). As SigLIP represents a single image with tokens and the downsampling ratio of the teacher tokenizer KL-16 (Rombach et al., 2022) is , we set the input resolution of the teacher tokenizer as to produce fine-grained tokens as denoising targets.
Instruction Tuning Data. When comparing with state-of-art LMMs in Table 4, our Ross is trained on approximately 1.2M instruction tuning data, which is a mixture of Cambrian-737K (Tong et al., 2024a) and SMR-473K (Zhang et al., 2024). Details of these two instruction tuning datasets are listed in Table 8 and Table 8, respectively. There might be some overlap but we simply concat these two datasets as it is already empirically effective.
Evaluation Prompts. We provide a thorough examination of all evaluation benchmarks utilized in this paper in Table 9. Notably, for MMVP (Tong et al., 2024b), which is not officially supported by VLMEvalKit (Duan et al., 2024), we follow Cambrian-1 (Tong et al., 2024a) to reformat the original question into a multiple-choice format and compute the accuracy using exact matching.
Computational Costs. As demonstrated in Table 10, the denoising process introduces a negligible increase in training time (10% compared to the baseline), while the benefits outweigh the minor additional costs.
Appendix C More Experiments
C.1 More Ablations
KL-16 v.s. VQ-16. Our default tokenizer is a continuous VAE (Kingma, 2013) with Kullback-Leibler (KL) divergence trained by Rombach et al. (2022). We further conduct experiments with a discrete tokenizer provided by Sun et al. (2024a), which is a VQGAN (Esser et al., 2021), i.e., VQVAE (Oord et al., 2017) with additional perceptual loss (Zhang et al., 2018) and adversarial loss (Goodfellow et al., 2014). KL-16 outperforms VQ-16. One intuitive explanation is that KL-16 preserves more low-level details than VQ-16 since quantization may lead to information loss. Moreover, quantitatively, on ImageNet (Deng et al., 2009) 256256 validation set, KL-16 achieves 0.87 rFID (Heusel et al., 2017) while the rFID (Heusel et al., 2017) of VQ-16 is 2.19.
Architecture of the Denoiser. Illustrated in Figure 10, the self-attention module is crucial, as original visual outputs are actually causal and we need to model inter-token discrepancy via self-attention. The number of trainable parameters is not the crux.
| Schedule of | POPE | Hallu. | MMVP | ChartQA | MMB |
| – | 87.9 | 55.0 | 29.3 | 34.0 | 73.8 |
| Linear (Ho et al., 2020) | 88.1 0.2 | 57.3 2.3 | 42.0 12.4 | 39.2 5.2 | 75.1 1.3 |
| Scaled Linear (Rombach et al., 2022) | 88.4 0.5 | 58.3 3.3 | 40.0 10.4 | 40.7 6.7 | 75.3 1.5 |
| GLIDE Softmax (Nichol et al., 2022) | 88.4 0.5 | 59.1 4.1 | 42.7 13.4 | 40.4 6.4 | 75.2 1.4 |
| GeoDiff Sigmoid (Xu et al., 2022) | 88.2 0.3 | 57.7 2.7 | 41.3 11.7 | 38.9 4.9 | 75.5 1.7 |
Schedule of . We study the effectiveness of different schedules of in Table 11. From the table, we can tell that even with different schedules of , Ross consistently improves the baseline, demonstrating its robustness to the denoising schedule.
Generative v.s. Reconstructive. We offer a detailed pipeline comparison in Figure 11. Experimental results have already been provided in Table 3. The implementation of generative methods is similar to Sun et al. (2024b) and Dong et al. (2024), where we adopt 576 learnable queries as inputs and take the corresponding outputs as conditions for the denoiser.
| CLIP-ViT-L/14@336 | SigLIP-ViT-SO400M/14@384 | |||||||
| Vicuna-7B-v1.5 | Qwen2-7B-Instruct | Vicuna-7B-v1.5 | Qwen2-7B-Instruct | |||||
| Benchmark | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross |
| POPE | 86.3 | 87.2 0.9 | 87.9 | 88.4 0.5 | 86.0 | 87.7 1.7 | 88.5 | 88.7 0.2 |
| HallusionBench | 52.5 | 55.8 3.3 | 55.0 | 59.1 4.1 | 50.4 | 53.8 3.4 | 57.3 | 58.2 0.9 |
| MMBench-EN | 67.0 | 67.6 0.6 | 73.8 | 75.2 1.4 | 64.5 | 69.2 4.7 | 76.3 | 76.9 0.6 |
| MMBench-CN | 60.0 | 59.8 0.2 | 72.9 | 73.7 0.8 | 63.1 | 63.4 0.3 | 75.7 | 76.3 0.7 |
| SEED | 66.7 | 66.4 0.3 | 70.3 | 70.7 0.4 | 68.2 | 69.0 0.8 | 72.3 | 72.1 0.2 |
| MMMU | 30.0 | 34.0 4.0 | 44.0 | 45.3 1.3 | 33.3 | 38.0 4.7 | 38.7 | 41.3 2.6 |
| MMMU | 35.3 | 36.0 0.7 | 41.9 | 42.6 0.7 | 34.2 | 35.4 1.2 | 41.8 | 43.8 2.0 |
| MMVP | 28.0 | 36.0 8.0 | 29.3 | 42.7 13.4 | 27.3 | 38.0 10.7 | 40.7 | 49.3 8.6 |
| AI2D | 61.2 | 61.4 0.2 | 71.9 | 73.3 1.4 | 62.6 | 62.4 0.2 | 74.0 | 74.5 0.5 |
| ChartQA | 32.9 | 39.8 6.9 | 36.2 | 41.6 5.4 | 34.0 | 48.2 14.2 | 44.4 | 46.9 2.5 |
| DocVQA | 33.4 | 41.6 8.2 | 31.1 | 44.7 13.6 | 40.4 | 40.7 0.3 | 39.2 | 39.3 0.1 |
| InfoVQA | 21.2 | 26.4 5.2 | 22.1 | 39.3 16.2 | 22.8 | 23.3 0.5 | 24.0 | 25.1 1.1 |
| TextVQA | 55.7 | 58.7 3.0 | 52.0 | 54.1 2.1 | 60.5 | 62.6 2.1 | 56.3 | 57.5 1.2 |
| OCRBench | 339 | 350 11 | 363 | 381 18 | 354 | 365 11 | 432 | 448 16 |
| RealWorldQA | 52.7 | 53.2 0.5 | 56.7 | 57.4 0.7 | 55.0 | 57.1 2.1 | 57.9 | 59.1 1.2 |
| Average | 47.8 | 50.6 2.8 | 52.1 | 56.4 4.3 | 49.2 | 52.4 3.2 | 55.4 | 56.9 1.5 |
We hypothesize that the underlying reason for the lower performance of generative methods in comprehension tasks is the weak correspondence between inputs and supervision under generative settings, which typically arises from both the (1) data and the (2) design of these methods.
(1) Typical generative methods that explore the synergy of comprehension and generation, usually leverage image generation conditioned on text instructions on (i) text-to-image datasets or (ii) interleaved datasets as extra supervision. However, (i) text-to-image datasets are typically designed to generate high-aesthetic samples rather than text-aligned ones, and (ii) interleaved datasets aim to enable few-shot learning via interleaving independent supervised examples, where reasoning becomes more important than alignment. Therefore, there exists a clear disconnect where the supervision (image) has little to do with the input (text instruction). For example, the CLIP-Score (Hessel et al., 2021), which measures the similarity between text and images, is only 0.3043 for the LAION-Art dataset (Schuhmann et al., 2022) and 0.2842 for the MMC4 dataset (Zhu et al., 2023), indicating that the supervision signals in these datasets are not well-suited for tasks requiring strong text-image alignment.
(2) Even when we attempt to ensure image-text alignment by converting aligned caption data into creation data for supervision, the results demonstrated in Table 3 remain unsatisfactory. This suggests that the design of generative objectives itself does not inherently require a strong correspondence between inputs and supervision targets.
In contrast, reconstructive methods like Ross leverage the original input images as auxiliary supervision, ensuring a strong and direct correspondence, which is crucial for tasks requiring accurate comprehension and interpretation of multimodal data, leading to significantly improved performance.
Extended Ablations on Different LLMs and Visual Encoders. We extend the ablation in Table 3 by incorporating more benchmarks, providing a more balanced and representative distribution of tasks. Empirical results in Table 12 demonstrate that our proposed vision-centric supervision utilized by Ross leads to significant improvements in most cases. Moreover, we found Ross contributes more significant improvements over fine-grained comprehension datasets, such as HallusionBench (Guan et al., 2024), MMVP (Tong et al., 2024b), ChartQA (Masry et al., 2022), and OCRBench (Liu et al., 2023c).
C.2 Comparison on High-Resolution Benchmarks
We incorporate the “anyres” technique proposed by LLaVA-v1.6 (Liu et al., 2024b) into our Ross. Specifically, for each image, we employ a grid configuration of 384{22, 1{2,3,4}, {2,3,4}1} to identify the input resolution, resulting in a maximum of 5729 3,645 visual tokens. Each 384384 crop is required to reconstruct the original input via the denoising objective proposed by Ross. In Table 13, our Ross-7B surpasses LLaVA-v1.6-7B (Liu et al., 2024b) and Cambrian-1-8B (Tong et al., 2024a) under most cases. These results indicate that Ross not only performs well at lower resolutions but also maintains its competitive edge at higher resolutions, making it a robust and versatile method.
| Model | ChartQA | DocVQA | InfoVQA | TextVQA | OCRBench | RealWorldQA |
| GPT-4V-1106 (OpenAI, 2023a) | 78.5 | 88.4 | – | 78.0 | 645 | 61.4 |
| Gemini-1.5 Pro (Team et al., 2023) | 81.3 | 86.5 | – | 78.1 | – | 67.5 |
| Grok-1.5 (xAI, 2024) | 76.1 | 85.6 | – | 78.1 | – | 68.7 |
| LLaVA-v1.5-7B‡ (Liu et al., 2024a) | 18.2 | 28.1 | 25.7 | 58.2 | 317 | 54.9 |
| LLaVA-v1.6-7B‡ (Liu et al., 2024b) | 65.5 | 74.4 | 37.1 | 64.8 | 532 | 57.6 |
| Cambrian-1-8B (Tong et al., 2024a) | 73.3 | 77.8 | – | 71.7 | 624 | 64.2 |
| Ross-7B | 76.9 | 81.8 | 50.5 | 72.2 | 607 | 66.2 |
| CLIP-ViT-L/14@336 | SigLIP-ViT-SO400M/14@384 | |||||||
| Vicuna-7B-v1.5 | Qwen2-7B-Instruct | Vicuna-7B-v1.5 | Qwen2-7B-Instruct | |||||
| Benchmark | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross |
| Vision-Language Benchmarks on Knowledge | ||||||||
| ScienceQA | 68.5 | 69.0 0.5 | 76.5 | 77.4 0.9 | 69.6 | 71.3 1.7 | 78.3 | 78.5 0.2 |
| MMMU | 30.0 | 34.0 4.0 | 44.0 | 45.3 1.3 | 33.3 | 38.0 4.7 | 38.7 | 41.3 2.6 |
| MMMU | 35.3 | 36.0 0.7 | 41.9 | 42.6 0.7 | 34.2 | 35.4 1.2 | 41.8 | 43.8 2.0 |
| AI2D | 61.2 | 61.4 0.2 | 71.9 | 73.3 1.4 | 62.6 | 62.4 0.2 | 74.0 | 74.5 0.5 |
| Language Benchmarks | ||||||||
| MMLU | 26.5 | 27.4 0.9 | 57.1 | 60.7 3.6 | 26.0 | 25.9 0.1 | 60.9 | 61.0 0.1 |
| HellaSwag | 27.0 | 26.9 0.1 | 46.4 | 46.2 0.2 | 27.1 | 27.0 0.1 | 45.5 | 46.6 1.1 |
| IFEval | 41.2 | 44.6 3.4 | 47.1 | 49.2 2.1 | 43.6 | 43.8 0.2 | 47.8 | 48.1 0.3 |
| IFEval | 28.7 | 35.3 6.7 | 35.1 | 37.0 1.9 | 32.5 | 33.1 0.6 | 35.3 | 36.2 0.9 |
| Average | 39.8 | 41.8 2.0 | 52.5 | 54.0 1.5 | 41.1 | 42.1 1.0 | 52.8 | 53.8 1.0 |
| Benchmark | 0.5B | 1.5B | 3B | 7B | ||||
| LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | |
| POPE | 50.0 | 60.4 10.4 | 85.3 | 87.9 2.4 | 87.3 | 88.1 0.8 | 87.9 | 88.4 0.5 |
| HallusionBench | 45.8 | 48.0 2.2 | 48.7 | 49.6 0.9 | 52.2 | 52.2 – 0.0 | 48.7 | 53.7 5.0 |
| MMBench-EN | 55.2 | 60.4 5.2 | 67.5 | 68.2 1.7 | 70.6 | 71.4 0.8 | 75.0 | 75.7 0.7 |
| MMBench-CN | 45.6 | 48.9 3.3 | 62.4 | 63.9 1.5 | 68.0 | 69.1 1.1 | 73.6 | 73.5 0.1 |
| SEED | 55.8 | 55.6 0.2 | 66.3 | 66.8 0.5 | 68.2 | 68.4 0.2 | 70.6 | 71.0 0.4 |
| OCRBench | 229 | 248 19 | 291 | 298 7 | 313 | 308 5 | 334 | 358 24 |
| MMMU | 35.2 | 36.0 0.8 | 44.7 | 45.0 0.3 | 48.7 | 49.0 0.3 | 48.0 | 48.0 – 0.0 |
| MMMU | 38.0 | 40.3 1.7 | 41.8 | 43.6 1.8 | 41.6 | 42.7 1.1 | 47.3 | 48.0 0.7 |
| AI2D | 45.3 | 46.0 0.7 | 59.0 | 59.5 0.5 | 62.9 | 63.2 0.3 | 68.3 | 68.5 0.2 |
| RealWorldQA | 45.1 | 46.4 1.3 | 50.5 | 53.5 3.0 | 55.7 | 57.9 2.2 | 59.5 | 59.9 0.4 |
| Average | 43.9 | 46.7 2.8 | 55.3 | 56.8 1.5 | 58.9 | 59.3 0.4 | 61.2 | 62.3 1.1 |
| PT | SFT | POPE | Hallu. | ChartQA | OCRBench | MMB | AI2D | |
| 558K | 737K | – | 87.9 | 55.0 | 34.0 | 363 | 73.8 | 72.4 |
| ✓ | 88.4 0.5 | 59.1 4.1 | 40.4 6.4 | 380 17 | 75.2 1.4 | 73.3 0.9 | ||
| 558K | 1.2M | – | 88.5 | 57.3 | 37.0 | 389 | 75.7 | 74.5 |
| ✓ | 88.8 0.3 | 57.8 0.5 | 42.0 5.0 | 392 3 | 76.8 1.1 | 74.7 0.2 | ||
| 2M | 737K | – | 88.1 | 55.6 | 37.3 | 384 | 76.2 | 72.3 |
| ✓ | 88.3 0.2 | 56.2 0.6 | 41.9 4.5 | 398 14 | 77.0 0.8 | 73.4 1.1 | ||
| 2M | 1.2M | – | 88.5 | 53.8 | 41.2 | 388 | 76.5 | 73.9 |
| ✓ | 88.9 0.4 | 57.3 2.5 | 43.2 2.0 | 405 17 | 78.0 1.5 | 74.1 0.2 |
C.3 More Analysis
Language Capabilities. One possible concern of Ross is that this type of low-level reconstruction may harm the high-level language capabilities. To investigate this issue, we evaluate multi-modal benchmarks that mainly require general knowledge following (Tong et al., 2024a), including ScienceQA (Saikh et al., 2022), MMMU (Yue et al., 2024), and AI2D (Hiippala et al., 2021). Furthermore, we incorporate representative language benchmarks, including general understanding on MMLU (Hendrycks et al., 2020) and HellaSwag (Zellers et al., 2019), and instruction-following on IFEval (Zhou et al., 2023). Empirical results in Table 14 demonstrate that Ross does not harm language capabilities as it brings improvements in most cases.
Model Scaling Properties. To study the stability and scalability of Ross across different model sizes, we use the Qwen2.5 series (Team, 2024) with varying sizes as the base language model while keeping the CLIP-ViT-L/14@336 (Radford et al., 2021) as the visual encoder. The pre-training data is LLaVA-558K (Liu et al., 2023a), and the instruction tuning data is LLaVA-665K (Liu et al., 2024a). The results, shown in Table 16, demonstrate that Ross brings improvements over the baseline (LLaVA) across different model sizes in most cases.
Data Scaling Properties. To study the impact of the training data scale, we used Qwen2-7B-Instruct (Yang et al., 2024a) as the base language model and CLIP-ViT-L/14@336 (Radford et al., 2021) as the visual encoder. We compared the performance of Ross and the baseline under different scales of training data. Table 16 demonstrates that Ross consistently brings significant improvements as the training data scale increases.
Gradient Analysis. To better explain the reasoning behind how the vison-centric supervision enables the model to focus on relevant areas of the image during VQA tasks, we provide qualitative comparison using GradCAM (Selvaraju et al., 2020) on MMVP (Tong et al., 2024b) in Figure 12, since GradCAM helps in understanding which parts of the image the model is focusing on, making the model’s decision-making process more transparent. In our analysis, we visualize the gradient of the second-to-last block of the LMM, regarding the option of the ground-truth answer as the target class. Specifically in this case, where the providing question is about the spider web, our proposed vision-centric supervision signals provide more reasonable gradients and urge LMMs to focus on relevant regions, i.e., the spider web, as the training evolves.
C.4 Qualitative Comparisons
We provide sufficient qualitative comparisons in Figure 13, Figure 14, Figure 15, and Figure 16 on HallusionBench (Guan et al., 2024), MMBench (Liu et al., 2023b) English dev split, MMVP (Tong et al., 2024b), and SpatialBench (Cai et al., 2024), respectively. In Figure 13, Figure 14, and Figure 15, we compare our Ross-7B with the instruction tuning baseline LLaVA-v1.5-7B (Liu et al., 2024a), the state-of-the-art open-source method using extrinsic assistance Cambrian-1-8B (Tong et al., 2024a), and GPT-4V (OpenAI, 2023a).
As demonstrated in Figure 13, where we highlight the wrong parts of each prediction in red, our Ross manages to correctly answer the question with reduced hallucinations even when GPT-4V fails. Cambrian-1 (Tong et al., 2024a) even fails to follow the instructions in the second example. This could be because a super huge SFT data (7M) may harm the instruction-following abilities of LMMs. Qualitative results shown in Figure 14 demonstrate both enhanced reasoning abilities (the first example), low-level comprehension capabilities (the second example), and spatial understanding skills (the third example). Figure 15 illustrates that our Ross is good at recognizing various visual patterns, implying that the introduced reconstructive vision-centric objective indeed makes up the visual shortcomings of the original visual encoder.
Appendix D Discussion
One limitation is that Ross does not have generation capabilities, since Ross is designed for enhanced multimodal comprehension, without the need to generate photorealistic aesthetic images. Furthermore, the gap in training data between comprehension and generation methods also matters. For instance, PixArt- (Chen et al., 2023a), which is one of the most efficient text-to-image models, was trained on nearly 400M images to model the pixel discrepancy just in the first training stage. By contrast, our Ross is only trained on nearly 3M images for one epoch. Future topics include achieving photorealistic text-to-image generation via incorporating more training samples.